# Herding Unmasked: Insights into Cryptocurrencies, Stocks and US ETFs

### v1

@@ Abstract
Herding behavior has become a familiar phenomenon to investors, carrying the potential danger of both undervaluing and overvaluing assets, while also threatening market stability. This study contributes to the literature on herding behavior by using a more recent dataset to cover the most impactful events of recent years. To our knowledge, this is the first study examining herding behavior across three different types of investment vehicle. Furthermore, this is also the first study observing herding at a community (subset) level. Specifically, we first explore this phenomenon in each separate type of investment vehicle, namely stocks, US ETFs and cryptocurrencies, using the widely recognized Cross Sectional Absolute Deviation (CSAD) model. We find similar herding patterns between stocks and US ETFs, while these traditional assets reveal a distinction from cryptocurrencies. Subsequently, the same experiment is implemented on a combination of all three investment vehicle types. For a deeper investigation, we adopt graph-based techniques such as Minimum Spanning Tree (MST) and Louvain community detection to partition the given combination into smaller subsets whose assets are most similar to each other, then seek to detect the herding behavior on each subset. We find that herding behavior exists at all times across all types of investment vehicle at a subset level, although the herding might not manifest at the superset level. Additionally, this herding behavior tends to stem from specific events that solely impact that subset of assets. Given these findings, investors can construct an appropriate investment strategy composed of their choice of investment vehicles they are interested in.
Keywords cryptocurrencies  $\cdot$ stocks  $\cdot$ US ETFs $\cdot$ herding behavior  $\cdot$ community detection
@@ 1 Introduction
Market participants have long been known to play a crucial role in shaping the price movements of financial assets [1]. High demand for an asset propels its prices upward, while a loss of interest or adverse news about it tends to lead to a price decrease. As a result, the understanding of investor reactions to various market conditions, e.g. financial crises, political events, bull and bear markets, is necessary to make wise investment decisions. Indeed, there have been a considerable number of studies focusing on herding behavior among investors in financial markets, from traditional ones like stocks and ETFs to alternative assets such as art real estate and crypto [2, 3, 4, 5, 6, 7, 8, 9]. Nowadays, this is still a topic of interest as financial markets are constantly changing and reacting to new events over time [2]. As defined by most studies, herding behavior describes collective actions from different investors, either purchasing or selling, targeting a set of different assets at a certain period which results in the same price movements between these assets [5]. This behavior typically arises from two scenarios [3]: First, intentional herding - where investors disregard their individual beliefs and follow the decisions made by the majority. Second, unintentional herding - where investors do not follow the market consensus and make a decision by themselves. However, in this case, they accidentally make the same decisions due to shared information sources such as financial reports, price momentum and analyses. It has been shown that the collective movements brought about by intentional and unintentional herding directly threaten market stability and are considered a key factor contributing to speculative bubbles [10].
@@
Different approaches have been proposed to detect herding signals. These can be classified into statistical and Neural Network methods. The former often involves constructing a regression model based on return values, then using optimization techniques such as Ordinary Least Squares (OLS) to obtain independent features’ coefficients, which are assessed to identify the herding signals, using methods like Cross-Sectional Standard Deviation (CSSD model) [11], Cross-sectional Absolute Deviation (CSAD model) [12], Power Law Exponent-based model [4] and return-variance based model [5]. An alternative is to construct a metric based on a database of investors (e.g. the number of investors and demographics data) and their trading activities (e.g. volume of purchasing and selling), then set a threshold for the metric to decide whether the herding appears. This approach needs more than the return values since a database of such trading activities is required, such as measures proposed by Lakonishok et al. [13], Frey et al. [6] and Patterson et al.[14]. Regarding Neural Networks, these methods are rather new in this area. Currently, one mainly deploys Deep Learning models such as Long Short Term Memory (LSTM) [15] and BERT [16] to gauge the market sentiment (e.g. positive, negative and neutral) which is then incorporated as an independent feature in regression models akin to statistical approaches [17]. To this end, statistical approaches remain predominant in herding detection studies, with the CSAD model emerging as the most commonly utilized option [18, 19, 20, 21, 22].
@@
One common finding agreed upon by most studies in different asset classes is that investors tend to follow the market consensus during periods of market uncertainty, which eventually results in herding behavior. Historical examples include events like Black Monday in 1987 [23], the Dot-Com/Tech crash in 2000 [23], the global financial crisis in 2007-2008 [24], and more recently the crash brought on by Covid-19 pandemic in 2020  [7]. Notably, it is worth mentioning that the role of algorithmic trading during market panics has been revealed, showing that this decision-making method still tends to operate during periods of high uncertainty [25]. However, this approach does not seem to reduce herding behavior. On the contrary, it appears to inadvertently cause herding due to identical decisions made by algorithmic trading systems [26]. In addition, studies in the cryptocurrency market indicate that its investors, who tend to be naïve, are more likely to behave irrationally and are easily influenced by price fluctuations as well as concrete items of news. By contrast, it has been suggested by a number of authors that professional investors tend to be dominant in traditional markets such as stocks, commodities and currencies. Besides, herding behavior in these markets is mainly explained by economic uncertainty as well as macroeconomic factors [8, 27, 4, 28].
@@
Given the diversification of studies in herding behavior detection, most of them only focus on one particular type of investment vehicle. Since investors tend to diversify their portfolios by investing in different types of investment vehicle [29], only considering one of them overlooks this investment manner, resulting in a shortage of insights. Our study addresses this limitation by using three popular types of investment vehicle, e.g. stocks, US ETFs and cryptocurrencies.
@@
Furthermore, existing literature tends to employ a herding detection model on the entire dataset comprising all available assets. As a result, conclusions regarding the herding phenomenon are obtained based solely on the whole dataset, potentially yielding inadequate results. For instance, the herding might just be true for a portion of the dataset while leaving the others unaffected but these latter are still counted because they belong to the same dataset. We propose an approach to tackle this limitation. In particular, we first use Louvain community detection [30] to group the assets based on their similarity in terms of price movements. Therefore, each group contains the assets that are most similar to each other on the basis of price movements. To recall, the fundamental idea of herding behavior is that assets display synchronized price movements due to the mutual trading actions of investors at the same time, especially during high market uncertainty. Thus, assets dissimilar to each other are less likely to demonstrate herding signals. To this end, our way of splitting assets can help reduce false positives in the results. Next, herding detection is applied to each group. Following this practice, we can observe the herding behavior at a community level, offering deeper insights compared to conducting herding detection on the entire dataset.
@@
To answer these questions, we utilize 30-min closing price time series of 222 assets, including 146 stocks, 49 US ETFs tracking their corresponding US indices and 27 cryptocurrencies. All of the data is sourced from the platform FirstRate Data. As mentioned earlier, the CSAD model is one of the most popular approaches for herding detection at the moment, thus, we also use this model in our present work for ease of comparison with others’ work. Regarding community detection, we first construct an asset graph based on the Pearson correlation [31] between the assets, then apply Minimum Spanning Tree (MST) [32] and Louvain [30] algorithms to the asset graph in order to identify the communities. The details of this process are described in Section 3.
Results obtained from our study can be summarised into the following core findings: Firstly, while herding patterns in stocks and US ETFs are similar, there are differences between these and cryptocurrencies, meaning that these digital assets still remain separate from traditional assets to some extent. Additionally, the herding signal is stronger in the US ETFs compared to stocks and cryptocurrencies (RQ1). Secondly, by detecting herding behavior at a subset (community) level, we find that there are always subsets (communities) revealing herding behavior in each type of investment vehicle throughout the period examined, even though this phenomenon is not pronounced in the corresponding superset at some stages. Therefore, we conclude that while herding may not manifest when looking at all assets, it can still be found within specific subsets (RQ2). Thirdly, we notice 4 stock sectors (i.e. Information Technology, Finance, Healthcare and Energy), wherein assets within each sector and their corresponding ETFs tend to form communities themselves. Furthermore, the herding behavior observed in these communities tends to stem from specific events that solely impact that sector (RQ2).
The remainder of the article is organized as follows: Section 2 goes through related works. Section 3 provides a description of the datasets followed by methodologies. Section 4 describes the experimental results followed by implications and hypothesis. Lastly, the conclusion of this study is given in Section 5.
@@ 2.1 Periods of Uncertainty in Recent Years
The US-China trade war took place in January 2018, following the imposition of initial tariffs on Chinese imports by the new US administration at that time [33]. The primary goals behind this decision were to reduce the US trade deficits, stop the requirements of American companies to transfer technology and protect US technology companies from intellectual property theft [36]. As a countermeasure, China also levied different tariffs on the US products [33]. Although the trade war started early in 2018, it escalated during 2019 [37]. Consequently, financial markets in both the US and China were significantly affected, especially in the commodity, energy and technology sectors [38, 39, 37]. Major stocks in both countries as well as the ETFs that track them fell sharply [40, 41]. Trade war headlines attracted huge attention throughout the year. The US President Donald Trump’s announcement relating to the tariff imposition caused the stock and index values to fluctuate dramatically [42]. Indeed, it has been shown that there was a wave of selling pressure in 2019 because the stock prices fluctuated due to the panic of investors caused by the trade war [43]. In China, the economy also exhibited a decline in GDP growth rate in 2019 [36], accompanied by decreased returns across various sectors in financial markets [44]. The uncertainty in financial markets due to this trade war was also seen in other countries across the world [45, 39]. The event ended in January 2020 after both countries signed off on the so-called "Phase One agreement" trade deal [46].
@@ 2.3 Herding Behavior in the Cryptocurrency Market
Similar to traditional markets, the Covid-19 pandemic has also attracted significant attention in the cryptocurrency market. In [7], the authors gathered intraday trading data for 9 large cryptocurrencies such as Bitcoin, Ethereum and Litecoin, covering a 2-year period from January 2019 to January 2021. They applied Patterson and Sharma’s herding intensity measure to each cryptocurrency in two distinct sub-periods: Pre-Covid-19 (January 2019 - December 2019) and Covid-19 pandemic (January 2020 - January 2021). Unlike other aforementioned approaches, this method can detect the herding activities in each asset. Results suggested a dramatically greater herding intensity during the latter sub-period across all examined cryptocurrencies. Nevertheless, while all cryptocurrencies showed the herding phenomenon during this time, no evidence of herding was found in Bitcoin, Ethereum, and Ripple during the Pre-Covid sub-period. These findings align with those of Youssef and Waked who executed a hybrid model combining CSAD and Markov-switching methods on 43 cryptocurrency price time series with daily frequency [21]. In particular, their dataset spans from 28/04/2013 to 11/11/2020, divided into two sub-samples based on the Covid-19 pandemic, including the Pre-Covid from the beginning until the end of 2019 and the Covid outbreak thereafter. As expected, only the second sample (i.e. the Covid outbreak) revealed the herding evidence. Furthermore, for the second part of the work, the authors used the coronavirus media coverage index (MCI) which calculates the percentage of all news sources covering the novel coronavirus topic to test the relationship between news and the herding magnitude. Interestingly, they found that the news about the pandemic significantly drove the herding behavior among investors in the cryptocurrency market, indicating the fact that cryptocurrency traders are mainly influenced by news and people’s sentiments. The result is in line with most of the existing literature [91, 92, 93, 94].
@@
More recently, a few studies have utilized a newer dataset that covers up to the year 2023 [95, 96]. However, instead of segmenting the sample into smaller intervals like the previous studies did, they mainly retained the whole period which caused ambiguity for herding detection models. Consequently, their findings tended to suggest an absence of herding throughout the examined period, even during highly uncertain times such as the outbreak of the Covid-19 pandemic.
From the existing literature, it is evident that most studies investigating the financial markets (both traditional and cryptocurrency) focus on periods before the year 2022, while only a few works extend into more recent times. To this end, our work contributes to the existing literature by using more up-to-date data, which spans from April 2019 until May 2023. Moreover, although several studies divided their data into smaller sub-periods, these divisions seem not to be sufficiently refined. To address this, we segment our data into finer sub-periods corresponding to major events that occurred in the financial markets throughout the examined period. Further details of this time division will be described in Section 3.

### v2

@@ Abstract
Herding behavior has become a familiar phenomenon to investors, with potential dangers of both undervaluing and overvaluing assets, while also threatening market stability. This study contributes to the literature on herding behavior by using a recent dataset, covering the most impactful events of recent years. To our knowledge, this is the first study examining herding behavior across three different types of investment vehicle and also the first study observing herding at a community (subset) level. Specifically, we first explore this phenomenon in each separate type of investment vehicle, namely stocks, US ETFs and cryptocurrencies, using the Cross-Sectional Absolute Deviation model. We find mostly similar herding patterns for stocks and US ETFs. Subsequently, the same experiment is implemented on a combination of all three investment vehicles. For a deeper investigation, we adopt graph-based techniques including the Minimum Spanning Tree and Louvain community detection to partition the combination into smaller subsets to detect herding behavior for each subset. We find that herding behavior exists at all times across all types of investment vehicle at a subset level, although perhaps not at the superset level, and that this herding behavior tends to stem from specific events that solely impact that subset of assets. Lastly, we explore herding by examining the financial contagion effects between these types of investment vehicle. Results show that US ETFs not only have a tendency to propagate similar trading behaviors in stocks and especially cryptocurrencies but also show self-reinforcing herding behavior, acting as drivers of their own trends.
@@ 1 Introduction
Market participants have long been known to play a crucial role in shaping the common price movements of financial assets [1]. High demand for an asset propels its prices upward, while a loss of interest or adverse news about the asset tends to lead to a price decrease. As a result, the understanding of investor reactions to various market conditions, e.g. financial crises, political events, bull and bear markets, is necessary to make wise investment decisions. Indeed, there have been a considerable number of studies focusing on herding behavior among investors in financial markets, from traditional ones like stocks and ETFs to alternative assets such as art real estate and cryptocurrency [2, 3, 4, 5, 6, 7, 8, 9]. This remains a topic of interest as financial markets are constantly evolving and reacting to new events over time [2]. As defined by most studies, herding behavior describes collective actions from different investors, either purchasing or selling, targeting a set of different assets at a certain period which results in similar price movements between these assets [5]. This behavior typically arises from two scenarios [3]: First, intentional herding - where investors disregard their individual beliefs and follow the decisions made by the majority. Second, unintentional herding - where investors do not follow the market consensus and make a decision by themselves. However, in this case, they accidentally make the same decisions due to shared information sources such as financial reports, price momentum and analyses. It has been shown that the collective movements brought about by intentional and unintentional herding directly threaten market stability and are considered a key factor contributing to speculative bubbles [10].
@@
Different approaches have been proposed to detect herding signals. These can be classified into statistical and Neural Network methods. The former often involves constructing a regression model based on return values, then using optimization strategies such as Ordinary Least Squares (OLS) to obtain independent features’ coefficients, which are assessed to identify the herding signals, using methods like Cross-Sectional Standard Deviation (CSSD model) [11], Cross-sectional Absolute Deviation (CSAD model) [12], the Power Law Exponent-based model [4] and the return-variance based model [5]. An alternative is to construct a metric based on a database of investors (e.g. the number of investors and demographics data) and their trading activities (e.g. volume of purchasing and selling), then set a threshold for the metric to decide whether the herding appears. This approach needs more than the return values since a database of such trading activities is required, such as measures proposed by Lakonishok et al. [13], Frey et al. [6] and Patterson et al.[14]. Regarding Neural Networks, these methods are rather new in this area. Currently, one mainly deploys Deep Learning models such as Long Short Term Memory (LSTM) [15] and BERT [16] to gauge the market sentiment (e.g. positive, negative and neutral) which is then incorporated as an independent feature in regression models akin to statistical approaches [17]. To this end, statistical approaches remain predominant in herding detection studies, with the CSAD model emerging as the most commonly utilized option [18, 19, 20, 21, 22].
@@
One finding common to most studies examining different asset classes is that investors tend to follow the market consensus during periods of economic uncertainty, which eventually results in herding behavior. Historical examples include events like Black Monday in 1987 [23], the Dot-Com/Tech crash in 2000 [23], the global financial crisis in 2007-2008 [24], and more recently the crash brought on by Covid-19 pandemic in 2020  [7]. Notably, it is worth mentioning that the role of algorithmic trading during market panics has been revealed, showing that this decision-making method still tends to operate during periods of high uncertainty [25]. However, this approach does not seem to reduce herding behavior. On the contrary, it appears to inadvertently cause herding due to identical decisions made by algorithmic trading systems [26]. In addition, studies in the cryptocurrency market indicate that its investors are more likely to behave irrationally and are easily influenced by price fluctuations as well as concrete items of news. In contrast, it has been suggested by a number of authors that professional investors tend to be dominant in traditional markets such as stocks, commodities and currencies. Besides, herding behavior in these markets is mainly explained by economic uncertainty as well as macroeconomic factors [8, 27, 4, 28].
@@
Given the diversification of studies in herding behavior detection, most of them only focus on one particular type of investment vehicle. Since investors tend to diversify their portfolios by investing in different asset types [29], only considering one of them overlooks this investment manner, resulting in a shortage of insights. Our study addresses this limitation by using three popular types of investment vehicle, e.g. stocks, US ETFs and cryptocurrencies.
@@
Furthermore, existing literature tends to employ a herding detection model on the entire dataset comprising all available assets. As a result, conclusions regarding the herding phenomenon are obtained based solely on the whole dataset, potentially yielding inadequate results. For instance, the herding might just be true for a portion of the dataset while leaving the others unaffected but the latter are still counted because they belong to the same dataset. We propose an approach to tackle this limitation. In particular, we first use Minimum Spanning Tree (MST) [30] and Louvain community detection [31] to group the assets based on their similarity in terms of price movements. Therefore, each group contains the assets that are most similar to each other. To recall, the fundamental idea of herding behavior is that assets display synchronized price movements due to the mutual trading actions of investors at the same time, especially during periods of high market uncertainty. Therefore, herding is more likely to occur within the assets that share similar price movements. By contrast, assets that are dissimilar to each other are less likely to demonstrate herding signals. Our way of clustering assets can help reduce false positives in the results since we assign assets with similar movement patterns to the herding detection models right at the beginning rather than a set of random assets with different price movements, which initially contradicts the fundamental nature of herding, as previously stated. Next, herding detection is applied to each group of similar assets. Following this practice, we can observe the herding behavior at a community level, offering clearer and more granular results compared to conducting herding detection on the entire dataset.
@@
To answer these questions, we utilize 30-min closing price time series of 222 assets, including 146 stocks, 49 US ETFs tracking their corresponding US indices and 27 cryptocurrencies. All of these data are sourced from the platform FirstRate Data. As mentioned earlier, the CSAD model is one of the most popular approaches for herding detection at the moment, thus, we also use this model in our present work for ease of comparison with others’ work. Regarding community detection, we first construct an asset graph based on the Pearson correlation [32] between the assets, then apply Minimum Spanning Tree (MST) [30] and Louvain [31] algorithms to the asset graph in order to identify the communities. The details of this process are described in Section Methodology and Data Collection. Lastly, and motivated by comments in the literature that herding and informational cascades on one asset result in similar effects on other assets [33], we investigate whether such contagion effects manifest themselves among the three investment vehicles in our study via a well-known approach Vector Autoregression (VAR), marking a further novelty in the paper.
Results obtained from our study can be summarised into the following core findings: Firstly, while herding patterns in stocks and US ETFs are mostly similar, there are differences between these and cryptocurrencies, meaning that these digital assets still remain separate from traditional assets in terms of herding behavior among their investors to some extent (RQ1). Secondly, by detecting herding behavior at a subset (community) level, we find that there are always subsets (communities) revealing herding behavior in each type of investment vehicle throughout the period examined, even though this phenomenon is not pronounced in the corresponding superset at some stages. Therefore, we conclude that while herding may not manifest when looking at all assets, it can still be found within specific subsets (RQ2). Thirdly, we notice 4 stock sectors (i.e. Information Technology, Financials, Healthcare and Energy), wherein assets within each sector and their corresponding ETFs tend to form communities themselves. Furthermore, the herding behavior observed in these communities tends to stem from specific events that solely impact that sector (RQ2). Fourthly, US ETFs reveal strong contagion effects to both stocks and cryptocurrencies, especially during times of market uncertainties, indicating their potential to be drivers of herding in financial markets. On the contrary, although stocks and cryptocurrencies also exhibit evidence of shock transmission to US ETFs, these effects are quite weak. Instead, US ETFs display an internal contagion phenomenon by transmitting shocks to themselves, potentially leading to a self-reinforcing herding behavior.
The remainder of the article is organized as follows: Section Related Works goes through related works. Section Methodology and Data Collection provides a description of the datasets followed by methodologies. Section Experimental Results and Discussion describes the experimental results followed by implications and hypothesis. Lastly, the conclusion of this study is given in Section Conclusions.
@@ 2.1 Periods of Uncertainty in Recent Years
The US-China trade war took place in January 2018, following the imposition of initial tariffs on Chinese imports by the new US administration at that time [34]. The primary goals behind this decision were to reduce the US trade deficits, stop the requirements of American companies to transfer technology and protect US technology companies from intellectual property theft [37]. As a countermeasure, China also levied different tariffs on the US products [34]. Although the trade war started early in 2018, it escalated during 2019 [38]. Consequently, financial markets in both the US and China were significantly affected, especially in the commodity, energy and technology sectors [39, 40, 38]. Major stocks in both countries as well as the ETFs that track them fell sharply [41, 42]. Trade war headlines attracted huge attention throughout the year. Announcements relating to the tariff imposition caused the stock and index values to fluctuate dramatically [43]. Indeed, it has been shown that there was a wave of selling pressure in 2019 because the stock prices fluctuated due to the panic of investors caused by the trade war [44]. In China, the economy also exhibited a decline in GDP growth rate in 2019 [37], accompanied by decreased returns across various sectors in financial markets [45]. The uncertainty in financial markets due to this trade war was also seen in other countries across the world [46, 40]. The event ended in January 2020 after both countries signed off on the so-called Phase One agreement trade deal [47].
@@ 2.3 Herding Behavior in the Cryptocurrency Market
Similar to traditional markets, the Covid-19 pandemic has also attracted significant attention in the cryptocurrency market. In [7], the authors gathered intraday trading data for 9 large cryptocurrencies such as Bitcoin, Ethereum and Litecoin, covering a 2-year period from January 2019 to January 2021. They applied Patterson and Sharma’s herding intensity measure to each cryptocurrency in two distinct sub-periods: Pre-Covid-19 (January 2019 - December 2019) and Covid-19 pandemic (January 2020 - January 2021). Unlike other aforementioned approaches, this method can detect the herding activities in each asset. Results suggested a dramatically greater herding intensity during the latter sub-period across all examined cryptocurrencies. Nevertheless, while all cryptocurrencies showed the herding phenomenon during this time, no evidence of herding was found in Bitcoin, Ethereum, and Ripple during the Pre-Covid-19 sub-period. These findings align with those of Youssef and Waked who executed a hybrid model combining CSAD and Markov-switching methods on 43 cryptocurrency price time series with daily frequency [21]. In particular, their dataset spans from 28/04/2013 to 11/11/2020, divided into two sub-samples based on the Covid-19 pandemic, including the Pre-Covid-19 from the beginning until the end of 2019 and the Covid-19 outbreak thereafter. As expected, only the second sample (i.e. the Covid-19 outbreak) revealed the herding evidence. Furthermore, for the second part of the work, the authors used the coronavirus media coverage index (MCI) which calculates the percentage of all news sources covering the novel coronavirus topic to test the relationship between news and the herding magnitude. Interestingly, they found that the news about the pandemic significantly drove the herding behavior among investors in the cryptocurrency market, indicating the fact that cryptocurrency traders are mainly influenced by news and people’s sentiments. The result is in line with most of the existing literature [93, 94, 95].
@@
More recently, a few studies have utilized a newer dataset that covers up to the year 2023 [96, 97]. However, instead of segmenting the sample into smaller intervals like the previous studies did, they mainly retained the whole period which caused ambiguity for herding detection models. Consequently, their findings tended to suggest an absence of herding throughout the examined period, even during highly uncertain times such as the outbreak of the Covid-19 pandemic. Additionally, most studies investigating the financial markets (both traditional and cryptocurrency) focus on periods before the year 2022. Given these current limitations, our work contributes to the existing literature by using more up-to-date data, which spans from April 2019 until May 2023. Moreover, we segment our data into finer sub-periods corresponding to major events that occurred in the financial markets throughout the examined period. Further details of this time division will be described in Section Methodology and Data Collection.
2.4 Contagion Effects in Financial Markets
In the study of financial contagion, understanding how shocks propagate across assets and markets is critical to assessing systemic risk and interdependencies. Financial contagion refers to the transmission of economic shocks between distinct markets or assets, which often leads to heightened volatility and amplified co-movement across financial instruments.Herding and Contagion are frequently considered together in the financial literature, especially in the context of cryptocurrencies [98], [99].Given evidence of herding within a set of assets, the primary transmitter of shocks can be identified through observed contagion effects. For instance, [100] examined herding behavior and contagion effects among 50 cryptocurrencies over a four-year period from 2015 to 2018. The authors found a weak herding signal within this group of cryptocurrencies, with Bitcoin being the main herding driver, as evidenced by its significant contagion effect on other coins.
Regarding contagion between asset classes, a number of authors have investigated such risks between cryptocurrencies and stock markets and a selection of these is described below. In [101], the topic of contagion from stock markets to Bitcoin (BTC) futures after BTC futures were launched is considered, indicating increased evidence of this after the launch. The authors further note that risk averse investors tend to avoid Bitcoin markets, preferring mature markets such as NASDAQ and NIKKEI, after the launch of BTC futures and in bear market periods. The authors in [102] compared cryptocurrency and stock market dynamics during Covid-19 and discussed co-movement and hidden regimes of cryptocurrencies and stocks in Covid-19 using the wavelet coherence and Markov switching approaches. The authors found that both markets correlate over time at low but not at high frequencies. Another study, [103] examined asymmetric contagion effects between stock index data and the largest two cryptocurrencies (ETH and BTC) using a time-varying symmetrized Joe-Clayton copula GARCH model with a nonlinear Granger causality test for data between 2017 and 2020. They found evidence of time-varying tail dependence, with lower tail dependence found to be greater relative to upper and asymmetric contagion between the stock and cryptocurrency markets. A recent paper [104] looks at contagion between cryptocurrencies and stocks in emerging and mature markets during the Covid-19 period with an approach using DCC Garch and Wavelet Cross-Correlation. The data featured BTC as a proxy for cryptocurrencies, MSCI on behalf of emerging markets and S&P500 closing prices from October 2014 to March 2022. Their findings show that during periods of financial upheaval (compared to more tranquil ones) an increase in conditional correlation was manifest. Evidence from wavelet cross-correlation analysis showed a positive cross-correlation between the BTC and equity markets in periods of financial unrest, with evidence for cross-correlation found at fine and coarse scales. Another recent paper dealing with a similar topic [105] found the cryptocurrency-stock correlation to be largely absent pre-March ’20 but increased considerably after. The data examined in the paper was unusual in that intraday as well as daily data were examined for BTC from July 2015 to June 2022. For stock indices during this period, SPY, DIA and QQQ were used as proxies for S&P500, Dow Jones, and Nasdaq, respectively. The authors speculated that the behavioral shift found was fueled by the Federal Reserve’s policy pandemic policy response to the Covid-19 pandemic. Also, little evidence of cryptocurrency shocks to stocks but large volatility spillovers in reverse were elicited. Finally, the authors linked increases in cryptocurrency-stock co-movement after the COVID-19 pandemic to the increased numbers of institutional investors in the cryptocurrency markets, where trades are sensitive to monetary policy changes.
In terms of contagion between stocks and ETFs, Da and Shive [106] saw that both information and non-fundamental shocks can be transmitted by ETFs. This takes place through an increased demand for ETFs resulting in price pressure, with onward transmission to the underlying basket of shares. The authors attribute this to arbitrageurs simultaneously taking opposite positions in the ETF and the underlying shares, with this effect being particularly noticeable for small and illiquid stocks. [107] took a network approach to examining risk spillover effects between ETFs and stocks. He looked at ETF-stock spillovers in different industrial sectors and also found - for certain sectors - a tendency for ETFs’ shocks to be transmitted to their constituent stocks. In addition, he commented on periods of high market volatility, when the contagion source tends to diversify. This effect is also noted in [108] who found, in addition, that ETFs also propagate tail event shocks from one stock to their other constituent stocks as well as propagate demand shocks. Gebka and Wohar [109] noted different price patterns in different sectors which they attributed to traders’ irrationality with this being particularly the case for basic materials, oil and gas and consumer services. The authors suggested that this phenomenon can be driven by investors moving in and out of markets in sync, an extreme preference for safe assets or overconfidence.